A consumer organization wants to estimate the actual tread wear index of a brand name of tires that claims graded 250 on thWill rate, thank you in advance.
Answer #1

Solution :

Given that,

sample size = n =23

Degrees of freedom = df = n - 1 = 23 - 1 = 22

\alpha=1-C%=1-.95=0.05

t\alpha /2,df = 2.074

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

= 2.074 * (20.7 / \sqrt 23)

=8.95

The 95% confidence interval estimate of the population mean is,

\bar x - E \leq \mu\leq\bar x + E

238.1-8.95 \leq \mu\leq 238.1+8.95

229.15 \leq \mu\leq 247.05

Ansb:

Yes. because a grade of 250 is not in interval

Ansc:

it is not unusual because it is just outside the confidence interval